OpenStudy (anonymous):

Which ordered pair is in the solution set of the following system of inequalities? y<1/2x+4 y>-x+1 a) (-5,3) b) (0,4) c) (3,-5) d) (4,0)

OpenStudy (anonymous):

You can either plug them in, or solve the system and see which points are in your solution set. Which would you prefer?

myininaya (myininaya):

(4,0) is the answer since 0<6=1/2*4+4 and 0>-3=-4+1 are both of the above inequalities are true statements